# KBAlign: Efficient Self Adaptation on Specific Knowledge Bases

### v3

@@ Abstract
Humans can utilize techniques to quickly acquire knowledge from specific materials in advance, such as creating self-assessment questions, enabling us to achieving related tasks more efficiently. In contrast, large language models (LLMs) usually relies on retrieval-augmented generation to exploit knowledge materials in an instant manner, or requires external signals such as human preference data and stronger LLM annotations to conduct knowledge adaptation. To unleash the self-learning potential of LLMs, we propose KBAlign, an approach designed for efficient adaptation to downstream tasks involving knowledge bases. Our method utilizes iterative training with self-annotated data such as Q&A pairs and revision suggestions, enabling the model to grasp the knowledge content efficiently. Experimental results on multiple datasets demonstrate the effectiveness of our approach, significantly boosting model performance in downstream tasks that require specific knowledge at a low cost. Notably, our approach achieves over 90% of the performance improvement that can be obtained by using GPT-4-turbo annotation, while relying entirely on self-supervision. We release our experimental data, models, and process analyses to the community for further exploration  111 https://github.com/thunlp/KBAlign.
KBAlign: Efficient Self Adaptation on Specific Knowledge Bases
@@ 1 Introduction
Figure 1: Comparison between different knowledge adaptation methods. KBAlign requires no external signals for tuning, and improves by itself in both tuning and inference stages.
Large language models (LLMs) have demonstrated their general capabilities across a wide range of downstream tasks (Achiam et al., 2023), while domain adaptation remains a crucial strategy to further enhance their performance (Ling et al., 2023). Although retrieval augmented generation (RAG) (Lewis et al., 2020) has become a common setting for LLMs to broaden the knowledge scope, tuning on specific domains can still help models outperform much larger general-purpose models (Zhao et al., 2024), highlighting the value of tailored training. Direct training on small-scale raw data, however, may degrade performance, necessitating alternative approaches to specific knowledge adaptation. Common solutions try to transform the data into more structured tasks like reading comprehension (Cheng et al., 2023), typically involve the incorporation of external knowledge signals (e.g., hand-crafted rules or GPT-4 annotations (Tan et al., 2024)). Nevertheless, it is actually proved that LLMs may achieve self improvement and refinement. For example, leveraging internal consistency and self-feedback (Liang et al., 2024), researchers have developed various filtering, correction, and preference training strategies to enhance the reliability of model outputs. Therefore, external knowledge signals is not necessary to some extent.
Considering that adaptation scenarios for specific knowledge could rapidly evolve and have some privacy restrictions, activating the potential of LLMs for autonomous adaptation is a better choice than adopting external annotations from human or online models. Drawing an analogy to human learning process, RAG is similar with open-book tests, in which students would conduct self study in advance. To be specific, they quickly read the fundamental content of the books by themselves, and can also verify their own answers. This may not be as good as spending time learning from external signals (e.g., related information from teachers), but is much more efficient. Therefore, the rapid self adaptation on specific knowledge materials is a useful capability.
@@
Inspire by the above analogy, we purpose KBAlign, a highly efficient self adaptation approach tailored to specific knowledge bases (KBs) comprising self-improvement learning. Generally, we tune the model in a completely self-supervised manner, and maximize the tuned model’s potential in inference. As shown in Fig. 1, for self annotation, we organize the original KB materials and conduct self annotation to get initial Q&A data pairs that can cover various downstream task scenarios. For iterative tuning, we require the model to check its own responses as the learning progresses, and help the model to modify the common mistakes at the current stage. Meanwhile, we conduct targeted inference in which strategies such as query expansion and confidence verification are adopted to refine the response.
@@
Experiments on fact QA, long-form QA, and professional field test have shown the effectiveness of our method across different backbone models. With a low cost, KBAlign masters the general knowledge content from the KB and achieves significant performance improvements. Side experiments including ablation studies and performance curve analyses identify the most efficient self-annotated data amount and optimal training volumes, offering valuable guidance for effectively applying KBAlign in practical scenarios.
@@
Our main contributions are as follows: (1) We propose KBAlign, a novel method for autonomous LLM adaptation tailored to knowledge bases. KBAlign is the first approach to help LLMs perform KB adaptation relying entirely on self annotations; (2) We provide empirical insights into efficient self adaptation to KBs, offering practical parameters and settings for deploying KBAlign; (3) We conduct a comprehensive analysis of the proposed self adaptation framework. Through a range of evaluation metrics and case studies, we identify the effectiveness of KBAlign and discuss the current limitations of our approach, highlighting areas for future improvement.
@@ 2 Related Work
Domain Adaptation.Though LLMs have shown their impressive capabilities across various scenarios (Jablonka et al., 2023), training methods for LLMs to adapt to certain domains still emerge in endlessly, due to the vertical application requirements. For domains with plenty of data resources, researchers usually directly take domain materials into pre-training (Wang et al., 2023a; Madani et al., 2023). In more cases, they continue to train based on general LLMs or mix the domain data with general corpus (Wu et al., 2023). To adopt the domain knowledge in a more efficient way, format conversion and annotation are often conducted (Zhang et al., 2024). For efficiency reasons, domain LLMs usually use parameter-efficient tuning to reduce the required data amount and training time (Ding et al., 2022). However, the construction of domain data is sometimes still costly. In our case, we mainly focus on the data perspective and try to propose more efficient data construction methods.
@@
Figure 2: Details for the KBAlign framework. (a) Schematic and data composition for the whole process; (b) Instances for different annotation strategies and tasks.
@@ 3.1 Task Setup
Current LLM adaptation approaches are usually used for domain transfer to enhance a type of capabilities (e.g., medicine, mathematics, coding). However, in real-world scenarios, there are often more specific needs, such as providing customized services based on user-specific document repositories, or plug-and-play integration of modules (e.g., operation manuals for new tools). These scenarios often have smaller data amount, and come with constraints such as confidentiality and convenience, leading to the inability to involve human annotation or rely on online large models, along with limited computational resources. In light of these constraints, RAG has emerged as a common approach to enhance the factual reliability of models. However, general retrieval methods often fail to leverage specific knowledge, therefore direct adaptation of models to specific KBs remains necessary.
We define KB adaptation as the process where, given a knowledge base $K$ (textual materials in our case), an original generative model $M$, and a retriever $R$, the goal is to efficiently enhance the grasp of the information contained within $K$, thereby improving performance on a test set $T$ based on $K$. Common forms of $T$ include knowledge question answering, multiple-choice tests, and other knowledge-intensive evaluations. The primary optimization objectives are to maximize performance scores on $T$ while minimizing training time within constrained resources.
There are two common approaches for leveraging $K$ to enhance model performance: training-based and inference-based methods. Training-based methods involve generating training data from $K$ using unsupervised techniques, such as masked language modeling or general language modeling on $M$, or targetly design $R$ for the current domain which is not covered in this work. Inference-based methods, on the other hand, focus on optimizing the retrieved content in our basic RAG setups, or post-processing the generated results to enhance relevance and accuracy. We will now introduce our method to combine the unsupervised training and the improved RAG, optimizing both the training and inference process. Examples for related process are shown in Fig. 2.
@@ 3.2 Self Annotation
To learn the knowledge from KBs without any supervised data, we conduct self annotation with the backbone model $M$ on the $K$ text. To be specific, we choose a paragraph of golden context $C_{g}$ as the knowledge source and require $M$ to raise s set of questions $Q$. We then supplement related context $C_{R}$ by the retriever $R$, and ask $M$ to annotate the answers $A$ based on the combined $C$. To increase the robustness and generalization, we provide a prompts set and randomly pick a style for the annotation, including concisely, in detail, and so on. $A$ is comparably precise because of the ensured existence of $C_{g}$ and our hand-crafted keyword filters (e.g., questions should not mention pronouns such as “in this paragraph”). $M$ is trained on <$Q$, $A$> instead of directly on $K$, which is similar with the self learning of human students converting and understanding the textbooks instead of directly reciting.
Owing to the diverse forms and attributes of $K$ and associated downstream tasks, we propose different organization strategies for $C$.
Short-dependency Annotation. For downstream tasks that prioritize precise fact knowledge expressed in one specific paragraph, we employ this approach to simply split $K$ into fixed-length chunks, each with no longer than $1,024$ words while keeps continuity of information across boundaries. One chunk is directly adopted as the annotation context $C_{g}$.
Long-Dependency Annotation. Considering that real-world tasks often require comprehensive understanding of multiple pieces of information at long distances in text, we design long-dependency annotation methods that split $K$ into shorter segments with less than $256$ words. Several segments $S_{1,...,n}$ with the same hierarchical directory, or with the highest embedding similarities across different directories, are concatenated together as $C_{g}$. When generating $Q$, the model is required to raise questions that: (1) involve knowledge from different segments to emphasize the multi-hop reasoning capability; (2) are as vague as possible, corresponding to a series of information $I_{1,...,n}$ annotated on $S$, based on which a refined long-form answer $A$ is generated to improve the integration capability.
@@ 3.3 Iterative Tuning
Apart from summarizing and self questioning to help deep understanding, human students also take tests at each learning stage and strengthen the knowledge they have not yet mastered by correcting their answers. Similarly, we hope that the model can improve itself through self-verifying in addition to understanding.
Initial Tuning. With the self-annotation <$Q$, $A$> data, we tune $M$ to get an initially adapted version. It is worth mentioning that a mixed paradigm has to be adopted to enhance the model generalization, that is, to provide a part of (50% in our practice) retrieved context $C_{R}$ which is different from the annotation $C$ as the model input combined with $Q$. This is to bridge the gap between tuning and inference in which RAG is not completely precise.
Self-Verify Tuning. Based on the $M_{0}$ that is initially adapted on a part of the data <$Q_{0}$, $A_{0}$>, we conduct RAG on $Q_{1}$ to get the predicted answer $P_{1}$ reflecting current capability of $M_{0}$. By providing the ground truth answer, the model verifies its own prediction and analyzes the wrong reason, which we name as $V_{1}$. In the next stage, we can then require the model to generate $V_{1}$ based on <$Q_{1}$, $P_{1}$>. And so on, we generate the verifying data based on current performance, and conduct the Q&A task and verify task at the same time in an iterative manner. In experiments, we use 25% of verify and 75% of Q&A, and implement 2-3 iterations.
@@ 3.4 Targeted Inference
We improve the downstream performance mainly by training the model to learn more specific knowledge. Practically, there are also strategies that can help optimize the generation results in inference stage. We employ Query Expansion (QE) to refine the retrieval results. To be specific, directly using the $Q$ as the search query may miss the useful information due to the short expression and the limitation of the retriever $R$. However, when our model has memorized the overall knowledge to a certain extent, it can read $Q$ and provide a prediction $P$ that is not necessarily accurate but very relevant to the KB content. We then expand the search query as $Q$+$P$, and this may help make the retrieval results much better.
The other strategy that can be used in reference is Self Verification, which is based on the verify capability learned in iterative verifying. For the generated $P$, the model can check the correctness by itself. It should be emphasized that this is not the standard strategy setting in subsequent experiments, because it will increase the time cost, and it is also difficult for the model to correct the error after realizing it. However, the model can at least provide an uncertainty warning, or sample a new response when the confidence score is low when needed, which helps improve reliability.
@@ 4.1 Datasets and Models
In order to evaluate the effectiveness of our method as comprehensively as possible, the following three datasets were used in the experiment:
LooGLE (Li et al., 2023). This is a long-text dataset, with textual materials that can be regarded as KBs and high-quality questions. We use the short-dependency data in LooGLE for retrofitting, combining altogether 2.2M tokens of text as $K$, and the corresponding 1,951 Q&A pairs for test. We evaluate the specific knowledge memorizing capability of the model in this dataset.
ASQA (Stelmakh et al., 2022). This is a long-form QA dataset. For each question, there exists several related segments of text from WikiPedia, and a comprehensive long answer that covers much information from them. We collect 794 Q&A pairs for test, their targeted segments and other related passages from WikiPedia, and get 1.8M tokens of text as $K$. We evaluate the knowledge recall and organizing capability of the model in ASQA test set, and do not use any training data from it.
JEC-QA (Zhong et al., 2020). This is a legal multiple choice dataset in Chinese. Related laws and reference books are seen as $K$, including 21M tokens of text. The train set has also not been used for adaptation, and the test set contains 1,985 of multiple choices. Through this set we evaluate the professional learning capability and instruction following in different inference formats.
@@
MiniCPM (Hu et al., 2024). This refers to MiniCPM-2.4B which is one of the backbone models in our experiment. It is an end-side LLM gaining the instruction following ability during pre-training, and has achieved the best performance among lightweight LLMs on several datasets. Therefore, we believe that it has wide personal applications and is suitable for efficient adaptation scenarios.
LLaMA-3.1. This refers to LLaMA-3.1-8B-Instruction  [^2] https://huggingface.co/meta-llama/Llama-3.1-8B which is aligned from one of the most popular open-source model families. We choose it to evaluate whether our method is universally helpful when the backbone model becomes stronger.
GPT-3.5. This refers to GPT-3.5-turbo-0125  [^3] https://platform.openai.com/docs/models which is a representative closed-source LLM with stable performance and comprehensive capability.

### v4

# KBAlign: Efficient Self Adaptation on Specific Textual Knowledge Bases

@@ Abstract
Although retrieval-augmented generation (RAG) remains essential for knowledge-based question answering (KBQA), current paradigms face critical challenges under specific domains. Existing methods struggle with targeted adaptation on small-scale KBs: vanilla unsupervised training exhibits poor effectiveness, while fine-tuning incurs prohibitive costs of external signals. We present KBAlign, a self-supervised framework that enhances RAG systems through efficient model adaptation. Our key insight is to leverage the model’s intrinsic capabilities for knowledge alignment through two innovative mechanisms: multi-grained self-annotation that captures global knowledge for data construction, and iterative tuning that accelerates convergence through self verification. This framework enables cost-effective model adaptation to specific textual KBs, without human supervision or external model assistance. Experiments demonstrate that KBAlign can achieve 90% of the performance gain obtained through GPT-4-supervised adaptation, while relying entirely on self-annotation of much smaller models. KBAlign significantly improves downstream QA accuracy across multiple domains with tiny costs, particularly benefiting scenarios requiring deep knowledge integration from specialized corpora. We release our experimental data, models, and process analyses to the community for further exploration 111 https://github.com/thunlp/KBAlign.
KBAlign: Efficient Self Adaptation on Specific Textual Knowledge Bases
@@ 1 Introduction
Figure 1: KBAlign schematic. We design special self-annotation methods to help master global KB knowledge, conduct iterative verifying to save training time costs, and adopt targeted inference to improve accuracy.
Large language models (LLMs) have demonstrated their general capabilities across a wide range of downstream tasks (Achiam et al., 2023), and the factual reliability of the models could be enhanced with common techniques such as retrieval augmented generation (RAG) (Lewis et al., 2020). When applied in specific domains, however, the adaptation of models to knowledge base (KB) materials remains a crucial strategy to further improve performance (Ling et al., 2023). Intuitively, adapted models align with the knowledge scope of the given KBs, rewriting the retrieval queries in a targetd manner and utilizing the KB information more fully. For example, when refering to the word “LLM”, both general QA model and the retriever may regard it as an AI term, while the aligned model can supplement keywords to the initial query according to the scope of legal KBs, and then generates a disambiguated response with the help of a more precise retrieved context, in which “LLM” stands for master of laws.
Existing adaptation methods usually construct domain models with large-scale training data (Zhao et al., 2024), while there are quite specific needs in real-world scenarios corresponding to small-scale textual KBs, such as providing customized services based on user-specific document repositories, or plug-and-play integration of modules. In these cases, simple LM training on raw data may degrade performance, necessitating alternative approaches to align with specific domains (Cheng et al., 2023). Targeted fine-tuning, on the other hand, typically involves the incorporation of external knowledge signals (Tan et al., 2024) to transform data into more structured tasks. When faced with constraints such as confidentiality, convenience, and limited computational resources, involving human annotation or relying on online large models becomes unpractictible. Therefore, a low-cost adaptation to small-scale KBs that does not rely on external supervisions is urgently needed.
@@
Drawing an analogy to the human learning process, RAG is similar to open-book tests in which students could query KB materials. If they conduct self-study in advance, quickly grasping the fundamental content of the books themselves, the effectiveness and efficiency in tests can be improved. Based on this idea, we purpose KBAlign, a highly efficient self-adaptation approach tailored to specific KBs comprising self-improvement learning. Generally, we align the model with the small domain in a highly efficient and completely self-supervised manner. As shown in Fig. 1, for self annotation, we organize the original KB materials in multiple grains and conduct self annotation to get instruction-response pairs that can cover various downstream task scenarios. For iterative tuning, we require the model to check its own responses and help modify common mistakes in the current stage for a faster convergence. Meanwhile, we conduct targeted inference in which strategies such as query expansion and confidence verification are adopted to refine the response.
@@
Experiments on fact QA, long-form QA, and professional field test have shown the effectiveness of our method across different backbone models. With a low cost, KBAligned models master the general knowledge content from the KB and achieves significant performance improvements. Side experiments including ablation studies and performance curve analyses identify the most efficient self-annotated data amount and optimal training volumes, offering valuable guidance for effectively applying KBAlign in practical scenarios.
@@
Our main contributions are as follows: (1) We propose KBAlign, a novel method for autonomous LLM adaptation tailored to textual KBs. It helps LLMs perform KB adaptation relying entirely on self annotations; (2) We provide empirical insights into efficient self-adaptation to KBs, offering practical parameters and settings for deploying KBAlign; (3) We conduct a comprehensive analysis of the proposed self-adaptation framework. Through a range of evaluation metrics and case studies, we identify the effectiveness of KBAlign and discuss the current limitations of our approach, highlighting areas for future improvement.
@@ 2 Related Work
Domain Adaptation.Though LLMs have shown their impressive capabilities in various scenarios (Jablonka et al., 2023), training methods for LLMs to adapt to certain domains still emerge in endlessly, due to the vertical application requirements. For domains with plenty of data resources, researchers directly take domain materials in pre-training (Wang et al., 2023a; Madani et al., 2023). In more cases, they continue to train based on general LLMs or mix domain data with the general corpus (Wu et al., 2023). To adopt domain knowledge in a more efficient way, format conversion and annotation are often performed (Zhang et al., 2024a) for fine-tuning. Some works focus on different settings for synthetic generation of QA data (Heydar et al., 2024; Ushio et al., 2023), while with the development of annotation model capabilities, the impact of specific synthetic strategies diminishes significantly. More crucially, existing approaches predominantly focus on local information and ignore global knowledge in synthesis.
@@
Figure 2: (a) Details for the KBAlign framework; (b) Instances for different annotation strategies and tasks.
@@ 3.1 Task Setup
We define KB adaptation as the process in which, given a knowledge base $K$ (textual materials in our case), an original generative model $M$, and a retriever $R$ for RAG, the goal is to efficiently align the models with the information in $K$ without any external signals, thus improving the knowledge-intensive tasks based on $K$. The optimization objectives are to maximize the performance scores in downstream tasks while minimizing training costs.
There are two common approaches utilizing $K$ to enhance model performance: tuning-based and inference-based methods. Tuning-based methods involve generating tuning data of $M$ from $K$ using unsupervised techniques, or designing specific $R$ for the current domain which is not covered in this work. Inference-based methods, on the other hand, focus on optimizing the retrieved content in the basic RAG setups, or post-processing the generated results to enhance relevance and accuracy. We now introduce our method which combines unsupervised tuning and RAG improvement, optimizing both the tuning and inference approaches. Examples are shown in Fig. 2.
@@ 3.2 Self Annotation
To learn the knowledge from KBs without any supervised data, we conduct self annotation with the backbone model $M$ on the $K$ text. To be specific, we choose a paragraph of golden context $C_{g}$ as the knowledge source and require $M$ to raise a set of questions $Q$. We then supplement the related context $C_{R}$ by the retriever $R$, and ask $M$ to annotate the answers $A$ based on $C_{g}+C_{R}$. When answering the questions with RAG, $M$ sometimes fails due to the vague context provided by $R$; while in the annotation process, $A$ is comparably precise because of the ensured existence of $C_{g}$ and our handcrafted keyword filters (e.g., questions should not mention pronouns such as “in this paragraph”).
Owing to the diverse forms and attributes of $K$ and associated downstream tasks, we propose multi-grained annotation corresponding to different organization strategies for $C$. The detailed process is shown in algorithm 1.
Short-dependency Annotation. For downstream tasks that prioritize precise fact knowledge expressed in one specific paragraph, we employ this approach to simply divide $K$ into fixed-length chunks, each with no more than $1,024$ words while keeping continuity of information across boundaries. One chunk is adopted directly as the annotation context $C_{g}$.
Long-Dependency Annotation. Considering that real-world tasks often require a comprehensive understanding of multiple pieces of information at long distances in text, we design long-dependency annotation methods that split $K$ into shorter segments with less than $256$ words. Several segments $S_{1,...,n}$ with the same hierarchical directory, or with the highest embedding similarities across different directories, are concatenated as $C_{g}$. When generating $Q$, the model is required to raise questions that: (1) involve knowledge from different segments to emphasize the multi-hop reasoning capability; (2) are as vague as possible, corresponding to a series of information $I_{1,...,n}$ annotated on $S$, based on which a refined long-form answer $A$ is generated to improve the integration capability.
@@ 3.3 Iterative Tuning
Apart from summarizing and self-questioning to help deep understanding, human students also take tests at each learning stage and strengthen the knowledge they have not yet mastered by correcting their answers. Similarly, we hope that the model can improve itself through self-verification in addition to understanding. The detailed process is also provided in algorithm 1.
Initial Tuning. With the self-annotation <$Q$, $A$> data, we tune $M$ to get an initially adapted version. Due to the limitations of the retriever, the retrieved $C_{R}$ in test scenario may differ from the annotation context $C$ (in which the golden paragraph $C_{g}$ must be included). Therefore, we randomly concatenate either $C$ or $C_{R}$ with the question $Q$ as the input. This mixed paradigm aims to bridge the gap between tuning and inference.
Self-Verify Tuning. We divide the annotated data into k parts <$Q_{1,2,...k}$, $A_{1,2,...k}$>, and adapt the model with the first part <$Q_{1}$, $A_{1}$> to get the initial version $M_{1}$. Using this model, we perform RAG inference in the second part $Q_{2}$ to obtain the predicted answer $P_{2}$ reflecting current capability of $M_{1}$. Given the ground-truth answer $A_{2}$, the model verifies its own prediction and analyzes the wrong reason, which we name $V_{2}$. In the next stage, we can then use <$Q_{2}$, $P_{2}$> as input and $V_{2}$ as output to continue tuning $M_{0}$. And so on, we generate the verification data based on current performance, and conduct the Q&A task and verify task at the same time in an iterative manner. In experiments, we use 25% of verify and 75% of Q&A, and implement 2-3 iterations.
@@ 3.4 Targeted Inference
We improve the downstream performance mainly by training the model to learn more specific knowledge. We also employ Query Expansion (QE) to refine the retrieval results in reference stage. To be specific, directly applying $Q$ as the search query may miss useful information due to the short expression and the limitation of the retriever $R$. Considering that our model has memorized the overall knowledge, it can provide a prediction $P$ that is relevant to the KB content. We then expand the search query as $Q$+$P$, and this may help make the retrieval results much better.
The other strategy that can be used in reference is Self Verification, which is based on the capabilities learned in iterative verifying. For the generated $P$, the model can check the correctness by itself. It should be emphasized that this is not the standard strategy setting in subsequent experiments, because it will increase the time cost, and it is also difficult for the model to correct the error after realizing it. However, the model can at least provide an uncertainty warning, or sample a new response when the confidence score is low when needed, which helps improve reliability.
@@ 4.1 Datasets and Models
In order to evaluate the effectiveness of our method as comprehensively as possible, we use four datasets in the experiment, each could form a corresponding KB (from 0.41 to 21 M tokens). Details are displayed in section A.
LooGLE (Li et al., 2023). This is a long-text dataset, with textual materials that can be regarded as KBs and high-quality questions. We evaluate the specific knowledge memorizing capability of the model in this dataset.
ASQA (Stelmakh et al., 2022). This is a long-form QA dataset. We evaluate the knowledge recall and organizing capability of the model in test set, and do not use any training data from it.
JEC-QA (Zhong et al., 2020). This is a legal multiple choice dataset in Chinese. We evaluate the professional learning capability and instruction following in different inference formats.
BioASQ  [^2] https://huggingface.co/datasets/kroshan/BioASQ. This is a biomedical question answering dataset. We evaluate the model’s biomedical knowledge retrieval and reasoning capabilities.
@@
MiniCPM (Hu et al., 2024). This refers to MiniCPM-2B which is one of the backbone models in our experiment. It is an end-side LLM gaining the instruction following ability during pre-training, and has achieved the best performance among lightweight LLMs on several datasets. Therefore, we believe that it has wide personal applications and is suitable for efficient adaptation scenarios.
LLaMA-3.1. This refers to LLaMA-3.1-8B-Instruct  [^3] https://huggingface.co/meta-llama/Llama-3.1-8B which is aligned from one of the most popular open-source model families. We choose it to evaluate whether our method is universally helpful when the backbone model becomes stronger.
GPT series. GPT-3.5 refers to GPT-3.5-turbo-0125  [^4] https://platform.openai.com/docs/models which is a representative closed-source LLM with stable performance and comprehensive capability. GPT-4o is an even stronger LLM.
LM. This represents directly conducting the language modeling task to align the model with KBs. Knowledgeable text is segmented into 512-token-length paragraphs, and mixed with general instruction tuning data (Ding et al., 2023) to keep the instruction following capability.
RAFT (Zhang et al., 2024b). This represents adapting language models to domain-specific RAG. Follow this method, we use GPT-4o to annotate data from the KBs, generating both Chain-of-Thought(CoT) reasoning and final answers to help the model focus on useful information while disregarding distractors. The annotated data used for supervised fine-tuning (SFT). We adopt this method as one of the baselines in our experiments.